You decide to test this screening test on a larger population, so you find 235501 people who agree to be tested and 5849 people test positive.

Given this information, you can determine the prevalence of disease in the population you studied. Report this in whole cases per 1000 persons per year.
Answer #1

Ans:

prevalence of disease=100*number of people with disease/total number of people

=100*5849/235501

=2.5 %

So, prevalence of disease( per 1000 persons per year)=(2.5/100)*1000=25 per 1000 person